OpenStudy (anonymous):

Which of the following functions re homomorphisms and why? a. f: Z to Z defined by f(x) = -x b. f: Z2 to Z2 defined by f(x)= -x c. g: Q to Q defined y g(x) = 1/ (x^2 + 1) d. h: R to M(R) defined by h(a) = [ (-a, 0) (a, 0)] e. f: Z12 to Z4 defined by f([x]12) = [x]4 where [u]n denotes the congruence class of the integer u in Zn.

terenzreignz (terenzreignz):

Group theory again :D You know what a homomorphism is?

OpenStudy (anonymous):

for rings of course you need to check two conditions, \[f(a+b)=f(a)+f(b)\] and \(f(ab)=f(a)f(b)\) so it is a little more stringent

OpenStudy (anonymous):

oh actually it did not say rings did it? i assumed that from the last problem are you viewing them as groups or rings?

OpenStudy (anonymous):

I'm sure they are rings

terenzreignz (terenzreignz):

You made me reconsider :/ Until the poster makes a specification, I'll be ambiguous :) Let's start with the first one, because I think it's the easiest. :D Let a and b be integers. Consider f(a+b) = -a-b = (-a) + (-b) = f(a) + f(b) Therefore, it's a GROUP homomorphism :) Let's check multiplication... consider f(ab) = -ab Hmmm.... Now check f(a)f(b) = (-a)(-b) = ab is ab = -ab ? Only when one of them is zero. If none of them is zero, then ab is not equal to -ab therefore f(ab) is not always equal to f(a)f(b) Therefore, it's NOT a RING homomorphism :P

OpenStudy (anonymous):

second is even easier in \(\mathbb{Z_2}\) there are only two elements, so \(f(x)=-x\) is the identity map

terenzreignz (terenzreignz):

Okay, @satellite73 second is easier :) But first is funner :D

terenzreignz (terenzreignz):

(more fun?) So, @urbanderivative There's your template for proving... take two arbitrary elements of the domain, and check if the mapping preserves both addition and multiplication :)

OpenStudy (anonymous):

off the top of my head i would say the third is not because a ring homo has to send 0 to 0

terenzreignz (terenzreignz):

Fourth sends a real number to a matrix?

OpenStudy (anonymous):

yeah, thats a matrix

terenzreignz (terenzreignz):

\[\huge h: \mathbb{R} \rightarrow M_\mathbb{R}\]\[\huge a \rightarrow \left[\begin{matrix}-a & 0 \\ a & 0\end{matrix}\right]\] this mapping?

OpenStudy (anonymous):

yes thats correct, I'm thinking it's a homomorphism, but I'm wondering if you can disprove it

terenzreignz (terenzreignz):

One sec... matrix multiplication makes me dizzy :D

OpenStudy (anonymous):

that's what I'm saying!

terenzreignz (terenzreignz):

Okay, clearly, it preserves addition, right?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Not convinced? :D Let a and b be real numbers. Then \[\huge f(a+b) = \left[\begin{matrix}-a-b & 0 \\ a+b & 0\end{matrix}\right]=\left[\begin{matrix}-a & 0 \\ a & 0\end{matrix}\right]+\left[\begin{matrix}-b & 0 \\ b & 0\end{matrix}\right]\]

OpenStudy (anonymous):

I'm pretty sure it* preserve multiplication, but I'm not sure on my matrix math

terenzreignz (terenzreignz):

:/ replace f by h...

terenzreignz (terenzreignz):

Okay, let's see... consider h(ab) first... \[\huge h(ab)=\left[\begin{matrix}-ab & 0 \\ ab & 0\end{matrix}\right]\]

terenzreignz (terenzreignz):

Now, let's have a look at h(a)h(b) \[\huge h(a)h(b)=\left[\begin{matrix}-a & 0 \\ a & 0\end{matrix}\right]\left[\begin{matrix}-b & 0 \\ b & 0\end{matrix}\right]\]

terenzreignz (terenzreignz):

And it's equal to \[\huge \left[\begin{matrix}ab & 0 \\ -ab & 0\end{matrix}\right]\]

terenzreignz (terenzreignz):

And once again, h(ab) = h(a)h(b) if and only if a = 0 or b = 0 any other case, they won't be equal Then, it wasn't a ring homomorphism after all :D

OpenStudy (anonymous):

Thank you! can you show me the equation for matrix math

terenzreignz (terenzreignz):

Equation for multiplication of matrices?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

I'm really not good with matrices, let alone explaining how the multiplication works, but here's the formula for 2x2 matrices :) \[\large \left[\begin{matrix}a & b \\ c & d\end{matrix}\right]\left[\begin{matrix}p & q \\ r & s\end{matrix}\right]=\left[\begin{matrix}ap+br & aq+bs \\ cp+dr & cq+ds\end{matrix}\right]\]